THE WAVELENGTH DEPENDENCE OF THE PHOTOLYSIS OF N-PENTALDEHYDE

\underline{L. ZHU} and T. J. CRONIN, Wadsworth Center, NYS Department of Health, Department of Environmental Health and Toxicology, State University of New York at Albany, Albany, NY 12201.

Aldehydes are pollutants formed from atmospheric photo-oxidation of hydrocarbons. Photodissociation of aldehydes represents a significant source of free radicals in the lower atmosphere. We have studied the uv photolysis of n-pentaldehyde in the 280-340 nm region by using dye laser photolysis in combination with cavity ring-down spectroscopy. Absorption cross sections of n-pentaldehyde were determined. The HCO radical was found to be a photodissociation product and its quantum yield as a function of photolysis wavelength will be presented. The dependence of HCO yield on n-pentaldehyde pressure and total pressure was examined. We have also investigated products formed from steady state photolysis of npentaldehyde at several incident wavelengths. The end-products were analyzed by using GC/MS and HPLC. Atmospheric implications of the results will be discussed.